# Thought2Text: Text Generation from EEG Signal using Large Language Models (LLMs)

### v1

@@ Abstract
Decoding and expressing brain activity in a comprehensible form is a challenging frontier in AI. This paper presents Thought2Text, which uses instruction-tuned Large Language Models (LLMs) fine-tuned with EEG data to achieve this goal. The approach involves three stages: (1) training an EEG encoder for visual feature extraction, (2) fine-tuning LLMs on image and text data, enabling multimodal description generation, and (3) further fine-tuning on EEG embeddings to generate text directly from EEG during inference. Experiments on a public EEG dataset collected for six subjects with image stimuli demonstrate the efficacy of multimodal LLMs (LLaMa-v3, Mistral-v0.3, Qwen2.5), validated using traditional language generation evaluation metrics, GPT-4 based assessments, and evaluations by human expert. This approach marks a significant advancement towards portable, low-cost "thoughts-to-text" technology with potential applications in both neuroscience and natural language processing (NLP).
@@ 1 Introduction
Brain-Computer Interface (BCI) systems, combined with portable and wearable noninvasive Electroencephalographic (EEG) devices, enable direct interfacing between the brain and external devices He et al. (2015). Advances in generating images and natural language using EEG Speier et al. (2016); Benchetrit et al. (2023); Défossez et al. (2023) hold promise for developing BCIs in various domains, including assistive communication (e.g., for ALS and stroke patients), mixed reality (AR/VR) experience enhancement, mental health diagnosis, and gaming. Recent strides in NLP driven by powerful LLMs such as OpenAI GPT-4. Achiam et al. (2023), Google Gemini Team et al. (2023), Meta-LLaMa Touvron et al. (2023), Mistral Jiang et al. (2023), and Microsoft Phi Gunasekar et al. (2023) have enabled multimodal integration, facilitating language generation from images Liu et al. (2024b) and speech Fathullah et al. (2024). Our research focuses on a multimodal solution to decode brain signals directly into text, leveraging EEG signals. We choose EEG for its affordability as compared to mainstream alternatives like functional Magnetic Resonance Imaging (fMRI) Tang et al. (2023), which are costlier and require complex setup.
Our approach for generating textual descriptions from EEG signals involves three key steps: (a) capturing language-agnostic EEG signals via visual stimuli, (b) encoding these signals into embeddings using a deep multichannel neural encoder, and (c) fine-tuning language models by projecting image and EEG embeddings into a token embedding space to generate responses. These responses are compared with gold standard image descriptions to compute the training loss. During inference, only EEG signals and a generic textual prompt are used as inputs to the LLMs to generate responses. Our method requires images, EEG data, and descriptions for training, while inference is bimodal, using only EEG to generate text.
For experiments, we used a public 128-channel EEG dataset from six participants viewing visual stimuli. Image descriptions were generated by GPT-4-Omni Achiam et al. (2023) and validated by experts, providing the necessary text modality for building EEG-to-text generation systems. Although the goal is text generation from EEG, we use a dataset with visual stimuli for their language-agnostic nature, avoiding cognitive biases linked to text reading (see Section 3 for details). We fine-tuned large language models (LLMs) using these descriptions, leveraging pre-trained instruction-based language models such as Mistral-v3 Jiang et al. (2023), LLaMa-v3 Touvron et al. (2023), and Qwen2.5 Bai et al. (2023a). Evaluation with standard generation metrics Sharma et al. (2017) and GPT-4-based assessments confirmed our approach’s effectiveness.
@@
Integration of brain signals with LLMs.
Fine-tuning models on EEG signals captured for visual stimuli, leveraging its language-agnostic nature to enhance LLM interaction and scalability across languages.
Validation of model efficacy on a popular public dataset that contains EEG signals captured using affordable devices in less constrained environments compared to traditional methods.
@@ 2 Related Work
Integrating behavioral signals like eye-gaze and brain signals into NLP and computer vision tasks Mishra and Bhattacharyya (2018); Sharma and Meena (2024) has seen significant progress. Key datasets include ZuCo 2.0 Hollenstein et al. (2019), which captures EEG and eye-gaze during natural language reading, and MOABB Jayaram and Barachant (2018), offering over 120,000 EEG samples from 400+ subjects from various BCI tasks such as motor imagery, visual evoked potentials, and cognitive load. Datasets such as CVPR2017 Spampinato et al. (2017) and MindBigData Vivancos and Cuesta (2022) provide substantial EEG data from responses to handwritten and open vocabulary object-based image stimuli. These datasets have facilitated research in classifying EEG data Spampinato et al. (2017); Palazzo et al. (2020); Khaleghi et al. (2023) and generating images from EEG signals using GANs Goodfellow et al. (2020) and latent diffusion models Rombach et al. (2022); Bai et al. (2023b); Lan et al. (2023); Tirupattur et al. (2018). However, using EEG signals in language generation remains elusive. The closest work to ours is EEG2TEXT Liu et al. (2024a), which uses EEG pretraining and a multi-view transformer to decode EEG signals into text. Unlike our approach, it does not leverage LLMs with prompt engineering for prompt-specific responses. We believe our method of fine-tuning LLMs with EEG input is the first of its kind.
@@
3 Dataset and the Need for Visual Stimuli
Building a system that generates text from thoughts naturally requires a dataset of paired examples. However, we believe that directly presenting participants with textual stimuli is not ideal, as using EEG data collected in this manner for thought analysis introduces additional complexities of language processing, such as determining brain activity windows for specific words, handling the retention of word context post-onset, and managing the overlap of contexts when words are shown in different time frames Wehbe et al. (2014); Murphy et al. (2022). Additionally, vocabulary size presents a challenge: while EEG-to-text systems perform well in closed-vocabulary settings, open-vocabulary decoding becomes inefficient as vocabulary size increases Martin et al. (2018); Wang and Ji (2022); Liu et al. (2024a). The core challenge in thought-to-text systems thus lies in two key aspects: (a) harnessing language-agnostic neural signals with minimal interference from linguistic processing and (b) using these signals to generate text in a target language, potentially by specifying an instruction or prompt.
This leads us to favor datasets where EEG signals are recorded in response to visual stimuli rather than textual input. For (a), visual stimuli provide several advantages: by relying on images, it circumvents the complexities of language processing and also elicits brain responses to salient image features, making them more suited for capturing neural activity in a language-agnostic manner. For (b), generating text from visual stimuli-evoked EEG remains challenging, as most EEG datasets collected with visual stimuli (e.g., CVPR2017 dataset Palazzo et al. (2020), MOABB Jayaram and Barachant (2018), MindBigData Vivancos and Cuesta (2022)) lack an associated text component, which is essential for evaluating whether the generated text from EEG signals accurately captures the perceived salient features of the images. To address this, robust image captioning tools like GPT-4 Omni Achiam et al. (2023) can be employed to generate captions for the visual stimuli, thereby creating a trimodal dataset with pairs. The text generation can further be guided by specifying a prompt or instruction to tailor the output to a particular language or context. This approach enriches the dataset and provides a more comprehensive foundation for thought-to-text research.
@@ 3 Dataset and the Need for Visual Stimuli
For our experiments, we utilize the CVPR2017 dataset Palazzo et al. (2020), which contains EEG data from six participants, each viewing 50 images across 40 diverse object categories (such as vehicles, musical instruments, etc.). Each EEG recording, corresponding to one participant and one visual stimulus, consists of 128 channels recorded for 0.5 seconds at a sampling rate of 1 kHz. This data is represented as a 128×N matrix, where N, approximately 500, represents the number of samples per channel in each segment. Since the exact duration of the EEG signals may vary, we follow the preprocessing approach of Palazzo et al. (2020) by discarding the first 20 samples (20 ms) to reduce interference from the previous image, and then standardizing the signal to a common length of 440 samples, accounting for segments with N < 500. The dataset provides pre-filtered signals across three frequency ranges: 14-70Hz, 5-95Hz, and 55-95Hz. In line with prior research, we select the 55-95 Hz range, as it has demonstrated the most reliable results.
@@
Since the CVPR2017 dataset lacks textual descriptions, we generate one-line captions for each image using GPT-4. We choose brief captions because the half-second EEG recordings may not capture enough information to support longer, more detailed descriptions. Given that these captions are auto-generated, it is crucial to ensure their quality and relevance. To validate the accuracy of the captions, we employed human annotators via Amazon Mechanical Turk (https://www.mturk.com/). These human experts reviewed the images and assessed the captions on a binary scale, judging both their fluency and adequacy. Table 1 summarizes the validation results, including the level of agreement between annotators when both rated a caption as acceptable, demonstrating the reliability of the auto-generated captions. We will release the descriptions under the same license as the original data.
| LLM | ROUGE-N | ROUGE-L | BLEU-N | METEOR | BERT | GPT-4 | GPT-4 | ||
| N=1 | N=2 | N=1 | N=4 | Score | Fluency | Adequacy | |||
@@
4 Method
Our approach employs a three-stage process to generate coherent text from EEG signals. The overall workflow is illustrated in Figure 1.
4.1 Stage1: Training EEG Encoder for Embedding Extraction
As shown in Figure 1, we utilize a multichannel EEG encoder inspired by ChannelNet Palazzo et al. (2020), a deep convolutional neural network model, to convert EEG signals into multidimensional embeddings (). These embeddings are further processed by an MLP classifier to predict object labels (), which correspond to objects present in the image stimuli. The label set matches Imagenet’s labels, available as part of the dataset, representing the most salient object in each image.
The model is trained by minimizing two losses: (A) a categorical cross-entropy loss () between the predicted and ground-truth object labels, and (B) a mean squared error () between the EEG embeddings () and pooled image embeddings () from a pretrained CLIP model Radford et al. (2021), which captures semantically rich image representations. CLIP’s embeddings offer robust transfer learning capabilities, making them ideal for aligning with EEG data.
To better align EEG embeddings with visual stimuli, we simplify images by removing non-central details like color, converting them into sketches using techniques such as Gaussian Blur and Canny filters. While this step is empirical and optional, it helps focus the alignment on core object features.
The overall loss function () balances and , weighted by a hyper-parameter (set to 0.5):
4.2 Stage2: Priming LLMs with Image Embeddings
To enable LLMs to process multimodal inputs, such as EEG and visual embeddings, we designed a projector inspired by recent advancements in vision-language models Zhang et al. (2024). Since LLMs are inherently text-based and cannot natively accept numerical embeddings, the projector transforms embeddings from the vision and EEG models ( in Stage 1 and in Stage 3) into the token embedding space of the LLM. This ensures the projected embeddings have the same dimensionality as the LLM’s token embeddings.
The projector is a simple feed-forward layer that maps the embeddings into the LLM’s token space. These transformed embeddings are then concatenated with token embeddings extracted from the input prompt, allowing the LLM to process both text and external modality embeddings seamlessly. In our main setting, the input prompt is structured as follows:
In stage 2, we first embed the tokens from the extended prompt, including the object labels, using the LLM’s token embedding layer. The token embeddings, , are then augmented with multimodal embeddings. Specifically, the embedding for the token is replaced by the projected multimodal embedding , and the token embedding for the is replaced by the embedding of the ground-truth object label. The multimodal embedding is computed by projecting external embeddings (e.g., from Stage 1) into the LLM’s token embedding space using the following transformation:
Here, and represent the parameters of the projector. Once is computed, it is prepended to the token embeddings from the input prompt, enabling the LLM to process multimodal information.
During training, labels are created by right-shifting the tokens in the prompt, aligning them with the ground-truth image descriptions. Special tokens such as beginning of sentences (), end of sentences () and padding tokens () are used considering LLM specific tokenizers. The input prompts are also converted into LLM specific chat templates. The LLM is then fine-tuned using a standard cross-entropy loss between the predicted tokens and the actual tokens from the ground-truth descriptions.
4.3 Stage3: Tuning LLMs with EEG Embeddings
This stage closely resembles Stage 2, with the distinction that we utilize , extracted from the EEG encoder trained in Stage 1, to compute the multimodal embeddings . During this stage, the parameters and of the projector undergo further training. We would like to highlight that throughout this and the previous stage, only the projector is trained, while the LLM and EEG encoders remain frozen to mitigate parameter instability caused by EEG noise.
4.4 Inference
During inference, the EEG encoder predicts the object label and EEG embeddings from raw EEG signals. The object labels are concatenated with a generic input prompt to compute token embeddings. The multimodal projector, trained in the previous stages, then computes from . These embeddings are concatenated and fed into LLMs to generate descriptions. Notably, no images are used during inference, making the inference process strictly bimodal.
| ID | Original Image Stimuli | Input | Ground Truth Object | Predicted Object (Stage 1) | Ground Truth Description (GPT-4) | Predicted Description (Thought2Text) |
| 1 |
EEG | mushroom | flower | A large yellow mushroom with a brown stem and a brown cap, surrounded by green foliage. | A group of mushrooms growing on a log. |
| 2 |
EEG + OBJ | piano | piano | A black grand piano in a living room. | A grand piano with a stool in front of it. |
| 3 |
EEG | piano | piano | A man in a red coat and black pants is playing a piano in a room with a chandelier. | A man is playing the piano in a dimly lit room. |
| 4 |
EEG + OBJ | pumpkin | pumpkin | A carved pumpkin with a face and eyes, sitting on a table. | A carved pumpkin with a spooky face on it. |
| 5 |
EEG + OBJ | flower | mushroom | A black and white photograph of a single daisy with a white center and a dark brown center. | A group of mushrooms growing on a log. |
| 6 |
EEG + OBJ | coffee mug | coffee mug | A hand holding a mug with a blue background and a handprint design. | A person holding a coffee mug with the words "World’s Best Dad" written on it. |
| 7 |
EEG | guitar | watch or watches | A young boy sitting on a chair playing a guitar. | A man is holding a guitar in front of a microphone. |
| 8 |
EEG + OBJ | camp or camping | camp or camping | A tent in a mountainous area with trees and fog. | A tent set up in a forest with a campfire nearby. |
5 Experimental Details
5.1 Dataset:
We utilize the open-source CVPR2017 dataset Palazzo et al. (2020) licensed for academic research, featuring EEG signals from six subjects viewing 50 images across 40 ImageNet classes Deng et al. (2009), totaling 2000 images. The data is split into training (), evaluation (), and test () examples. For more details, see Section 3.
5.2 Model Details
We evaluate three LLMs: LLaMa-v3 (meta-llama/Meta-Llama-3-8B-Instruct), Mistral-v03 (mistralai/Mistral-7B-Instruct-v0.3), and Qwen2.57B (Qwen/Qwen2.5-7B-Instruct), selected for their efficiency on consumer-grade GPUs like the NVIDIA RTX 3090. The multimodal embedding is projected to match the token embedding dimensions required by each LLM. All models are permissively licensed for academic research, and training takes approximately 7 GPU hours per LLM training cycle. During inference, we use a batch size of 1, with generation parameters such as top_k, top_p, and temperature set to their default values for each LLM.
5.3 Evaluation:
We use standard NLG metrics such as BLEU, METEOR, ROUGE Sharma et al. (2017), and BERTScore Zhang et al. (2019). Additionally, we assess generation quality using GPT-4, following Liu et al. (2024b).GPT-4 measures two aspects: fluency, for grammar , and adequacy, for accuracy in conveying meaning, both rated on a scale of 1-5, with 5 denoting highest quality.

### v2

@@ Abstract
Decoding and expressing brain activity in a comprehensible form is a challenging frontier in AI. This paper presents Thought2Text, which uses instruction-tuned Large Language Models (LLMs) fine-tuned with EEG data to achieve this goal. The approach involves three stages: (1) training an EEG encoder for visual feature extraction, (2) fine-tuning LLMs on image and text data, enabling multimodal description generation, and (3) further fine-tuning on EEG embeddings to generate text directly from EEG during inference. Experiments on a public EEG dataset collected for six subjects with image stimuli and text captions demonstrate the efficacy of multimodal LLMs (LLaMA-v3, Mistral-v0.3, Qwen2.5), validated using traditional language generation evaluation metrics, as well as fluency and adequacy measures. This approach marks a significant advancement towards portable, low-cost "thoughts-to-text" technology with potential applications in both neuroscience and natural language processing.
@@ 1 Introduction
Brain-Computer Interface (BCI) systems, combined with portable and wearable noninvasive Electroencephalographic (EEG) devices, enable direct interfacing between the brain and external devices He et al. (2015). Advances in generating images and natural language using EEG Speier et al. (2016); Benchetrit et al. (2023); Défossez et al. (2023) hold promise for developing BCIs in various domains, including assistive communication (e.g., for ALS and stroke patients), mixed reality (AR/VR) experience enhancement, mental health diagnosis, and gaming. Recent strides in NLP driven by powerful Large Language Models (LLMs) such as OpenAI GPT-4. Achiam et al. (2023), Google Gemini Team et al. (2023), Meta-LLaMA Touvron et al. (2023), Mistral Jiang et al. (2023), and Microsoft Phi Gunasekar et al. (2023) have enabled multi-modal integration, facilitating language generation from images Liu et al. (2024b) and speech Fathullah et al. (2024). Our research focuses on a multi-modal solution to decode brain signals directly into text, using EEG signals. We choose EEG because of its affordability compared to mainstream alternatives such as Functional Magnetic Resonance Imaging (fMRI) Tang et al. (2023), which are costlier and require complex setup. For generating text, we leverage LLMs, which enable flexible, high-quality text generation across various modalities, such as images Liu et al. (2024b), audio Rubenstein et al. (2023); Fathullah et al. (2024), and more and often outperform current open vocabulary task-specific methods Shi et al. (2024).
Our approach for generating textual descriptions from EEG signals involves three key steps: (a) capturing language-agnostic EEG signals via visual stimuli, (b) encoding these signals into embeddings using a deep multichannel neural encoder, and (c) fine-tuning language models by projecting image and EEG embeddings into a token embedding space to generate responses. These responses are compared with gold standard image descriptions to compute the training loss. During inference, only EEG signals and a generic textual prompt are used as inputs to the LLMs to generate responses. Our method requires images, EEG data and descriptions for training, while inference is bimodal, using only EEG to generate text.
For experiments, we use a public 128-channel EEG dataset from six participants viewing visual stimuli. The image descriptions are generated by GPT-4-Omni Achiam et al. (2023) and quality-checked by human annotators, providing the text modality necessary to build EEG-to-text generation systems. Although the goal is the generation of text from EEG, we use a dataset with visual stimuli for their language-agnostic nature, avoiding the potential complexities associated with reading and language processing (see Section 3 for details). We fine-tune large language models (LLMs) using these descriptions, leveraging pre-trained instruction-based language models such as Mistral-v3 Jiang et al. (2023), LLaMA-v3 Touvron et al. (2023), and Qwen2.5 Bai et al. (2023a). Evaluation with standard generation metrics Sharma et al. (2017) and GPT-4-based assessments confirmed the effectiveness of our approach.
@@
Integration of brain signals with instruction-tuned LLMs.
Fine-tuning models on EEG signals captured for visual stimuli, leveraging its language-agnostic nature to enhance LLM interaction.
Validation of the efficacy of the model in a popular public dataset that contains EEG signals captured using affordable devices.
@@ 2 Related Work
Integrating behavioral signals such as eye movement and brain signals into NLP and computer vision tasks Mishra and Bhattacharyya (2018); Sharma and Meena (2024) has seen significant progress. Key datasets include ZuCo 2.0 Hollenstein et al. (2019), which captures EEG and eye-gaze during natural language reading, and MOABB Jayaram and Barachant (2018), offering over 120,000 EEG samples from 400+ subjects from various BCI tasks such as motor imagery, visual evoked potentials, and cognitive load. Datasets such as MindBigData Vivancos and Cuesta (2022) and CVPR2017 Spampinato et al. (2017) provide substantial EEG data collected from participants’ responses to handwritten and open vocabulary object-based image stimuli respectively. These datasets have facilitated research on the classification of EEG data Spampinato et al. (2017); Palazzo et al. (2020); Khaleghi et al. (2023) and the generation of images from EEG signals using GANs Goodfellow et al. (2020) and latent diffusion models Rombach et al. (2022); Bai et al. (2023b); Lan et al. (2023); Tirupattur et al. (2018).
Additionally, multimodal datasets like The Alice Dataset Bhattasali et al. (2020), which includes EEG and fMRI recordings from participants listening to a story, provide two measurable modalities: audio stimulus and the corresponding text. Another recent multimodal dataset, EIT-1M Zheng et al. (2024), contains one million EEG-Image-Text pairs, collected as participants viewed visual-textual stimuli. At the time of writing, a partial version of the EIT-1M dataset was released, containing data for only one subject.
Generating language from EEG signals remains an elusive challenge. The most closely related work is EEG2TEXT Liu et al. (2024a), which utilizes EEG pretraining and a multi-view transformer to decode EEG signals into text. Another approach similar to ours, using multiple modalities, is presented in Ikegawa et al. (2024), where intracranial EEG (iEEG) signals were recorded from patients watching videos, and each video frame was used to generate images and text using CLIP vision-language model Radford et al. (2021). Unlike our approach, this study involved implanting electrodes in patients. Furthermore, these works do not leverage large language models (LLMs) with prompt engineering for generating prompt-specific responses. We believe our method of fine-tuning LLMs using non-invasive EEG input is the first of its kind.
3 Dataset and the need for Visual Stimuli
Building a system that generates text from neural activity naturally requires a dataset of paired examples. However, using textual stimuli presents inherent challenges. Reading is a learned skill that requires the decoding of symbols into sounds and meanings, syntactic parsing, and sequential integration in time. In contrast, visual perception is more innate, natural, and image processing is more parallel Dehaene (2009); Townsend (1990). Furthermore, using EEG data collected on textual stimuli introduces additional complexities of language processing, such as determining brain activity windows for specific words, managing retention of word context post-onset, and managing the overlap of contexts when words are shown in different time frames Wehbe et al. (2014); Murphy et al. (2022). Additionally, vocabulary size presents a challenge: while EEG-to-text systems perform well in closed-vocabulary settings, open-vocabulary decoding becomes inefficient as vocabulary size increases Martin et al. (2018); Wang and Ji (2022); Liu et al. (2024a). The core challenge in thought-to-text systems thus lies in two key aspects: (a) harnessing language-agnostic neural signals with minimal interference from linguistic processing and (b) using these signals to generate text in a target language, potentially by specifying an instruction or prompt.
This leads us to experiment with datasets where EEG signals are recorded in response to visual stimuli rather than textual input. For (a), visual stimuli provide several advantages: By relying on images, it circumvents the complexities of language processing and also elicits brain responses to salient image features, making them more suited to capture neural activity in a language-agnostic manner. For (b), generating text from visual stimuli-evoked EEG remains challenging, as most EEG datasets collected with visual stimuli (e.g., CVPR2017 dataset Spampinato et al. (2017), MOABB Jayaram and Barachant (2018), MindBigData Vivancos and Cuesta (2022)) lack an associated text component, which is essential for evaluating whether the generated text from EEG signals accurately captures the perceived salient features of the images. To address this, robust image captioning tools such as GPT-4 Omni Achiam et al. (2023) can be employed to generate captions that effectively capture the salient imformation of visual stimuli, thus creating a tri-modal dataset with tuples. Text generation can further be guided by specifying a prompt or instruction to tailor the output to a particular language or context. This approach enriches the dataset and provides a more comprehensive foundation for thought-to-text research.
@@ 3 Dataset and the need for Visual Stimuli
For our experiments, we utilize the CVPR2017 dataset Spampinato et al. (2017), which contains preprocessed EEG data from six participants, each viewing 50 images across 40 diverse object categories111we use the terms object and class interchangeably to refer to an object category. (such as vehicles, musical instruments, etc). Each EEG recording, corresponding to one participant and one visual stimulus, consists of 128 channels recorded for 0.5 seconds at a sampling rate of 1 kHz. Data are represented as a 128×N matrix, where N, approximately 500, represents the number of samples per channel in each segment. According to Spampinato et al. (2017), the EEG signals were pre-processed by first applying a second-order Butterworth bandpass filter between 5 Hz and 95 Hz and a notch filter at 50 Hz to remove power line noise. In addition, since the exact duration of the EEG signals can vary, the first 20 samples (20 ms) were discarded to reduce interference from the previous image, and then standardizing the signal to a common length of 440 samples, accounting for segments with N < 500. The dataset provides pre-filtered signals across three frequency ranges: 14-70Hz, 5-95Hz, and 55-95Hz. In line with previous research Palazzo et al. (2020), we selected the 55-95 Hz range, as it has shown the most reliable results. We used the training, validation, and test split from the original paper for overall and subject-wise analysis.
@@
Since the CVPR2017 dataset lacks textual descriptions, we used GPT-4 to generate brief one-line captions for each image as they are efficient in capturing salient information without introducing extraneous details. Here, the aim is to map EEG data to concise captions to simplify the alignment task, making it feasible despite the noise and variability inherent in EEG signals. To ensure quality, we validated these auto-generated captions using human annotators from Amazon Mechanical Turk, who rated their fluency and adequacy on a binary scale. Table 1 shows the validation results, including annotator agreement on acceptable captions, highlighting the reliability of the generated captions. We release the descriptions under the same license as the original data.
4 Method
Our approach uses a three-stage process to generate coherent text from EEG signals. The overall workflow is illustrated in Figure 1.
4.1 Stage1: Training EEG Encoder for Embedding Extraction
As shown in Figure 1, we use a multichannel EEG encoder inspired by ChannelNet Palazzo et al. (2020), a deep convolutional neural network model, to convert EEG signals into multidimensional embeddings (). These embeddings are further processed by an MLP classifier to predict object labels (), which correspond to objects present in the image stimuli. The label-set matches ImageNet’s labels, available as part of the dataset, representing the most salient object in each image.
The model is trained by minimizing two losses: (A) a categorical cross-entropy loss () between the predicted and ground-truth object labels, and (B) a mean squared error () between the EEG embeddings () and pooled image embeddings () from a pretrained CLIP model Radford et al. (2021), which captures semantically rich image representations. CLIP’s embeddings offer robust transfer learning capabilities, making them ideal for aligning with EEG data. Performance metrics for EEG-to-image classification using trained MLP Classifier
To better align EEG embeddings with visual stimuli, we simplify images by removing non-central details like color, converting them into sketches using techniques such as Gaussian Blur and Canny filters. Although this step is empirical and optional, using sketches help focus the alignment on the core features of the object as discussed in Fine-Grained Sketch-Based Image Retrieval Luo et al. (2023) and Interactive Sketch Question Answering Lei et al. (2024).
The general loss function () balances and , weighted by a hyper-parameter (set to 0.5):
4.2 Stage2: Priming LLMs with Image Embeddings
To enable LLMs to process multimodal inputs, such as EEG and visual embeddings, we designed a projector inspired by recent advancements in vision-language models Zhang et al. (2024) and is a fully connected feed-forward layer. Since LLMs are inherently text-based and cannot natively accept non-text embeddings, the projector transforms embeddings from the vision and EEG models into the token embedding space of the LLM. This ensures that the projected embeddings have the same dimensionality as the LLM’s token embeddings.
The projector is a simple feed-forward layer that maps the embeddings to the LLM token space. These transformed embeddings are then concatenated with token embeddings extracted from the input prompt, allowing the LLM to process both text and external modality embeddings seamlessly. In our main setting, the input prompt is structured as follows:
In stage 2, we first embed the tokens from the extended prompt, including the object labels, using the LLM’s token embedding layer. For this, text tokens are input as token IDs that were converted to dense vectors via a lookup in the LLM’s embed_tokens layer. The token embeddings, , are then augmented with multimodal embeddings. Specifically, the embedding for the token is replaced by the projected multimodal embedding , and the token-embeddings for the is replaced by the embeddings of the ground-truth object label. The multimodal embedding is computed by projecting pooled embeddings from CLIP, , into the LLM’s token embedding space using the following transformation:
Here, and represent the projector parameters. Once is computed, it is prepended to the token embeddings from the input prompt, enabling the LLM to process multimodal information.
During training, labels are created by right-shifting the tokens in the prompt, aligning them with the ground-truth image descriptions. Special tokens such as beginning of sentences (), end of sentences () padding tokens (), system and user and assistant message indicator tokens are used considering LLM specific tokenizers. The input prompts are also converted into LLM specific chat templates. The LLM is then fine-tuned using a standard cross-entropy loss between the predicted tokens and the actual tokens from the ground-truth descriptions.
4.3 Stage3: Tuning LLMs with EEG Embeddings
This stage closely resembles stage 2, with the distinction that instead of , we utilize , extracted from the EEG encoder trained in stage 1, to compute multimodal embeddings . During this stage, the projector parameters and are further tuned. We would like to highlight that throughout this and the previous stage, only the projector is trained, while the LLM and EEG encoders remain frozen to mitigate parameter instability caused by EEG noise.
4.4 Inference
During inference, the EEG ChannelNet encoder processes EEG signals to generate EEG embeddings . The multimodal projector, trained in stage 3, transforms into multimodal embeddings . is also fed into the MLP Classifier, trained in stage 1, to predict the object label. This predicted label is appended to a generic input prompt, similar to the one mentioned in Section 4.2, and the token embeddings are computed for the combined input. Finally, the token embeddings from the projector and the language prompt are concatenated and fed into the LLM to generate descriptions. Notably, no images are used during inference, making the process strictly bimodal.
5 Experimental Details
In this section we highlight the dataset, model details, and evaluation procedure.
5.1 Dataset
We utilize the open-source CVPR2017 dataset Spampinato et al. (2017), licensed for academic research, featuring EEG signals from six subjects viewing 50 images across 40 ImageNet classes Deng et al. (2009), totaling 2000 images. The data is split into training (), evaluation (), and test () examples. Additional details can be found under Section 3.
5.2 Model Details
We evaluate three LLMs: LLaMA-v3 (meta-llama/Meta-Llama-3-8B-Instruct), Mistral-v0.3 (mistralai/Mistral-7B-Instruct-v0.3), and Qwen2.5-7B (Qwen/Qwen2.5-7B-Instruct), selected for their efficiency on consumer-grade GPUs such as the NVIDIA RTX 4060Ti. The multimodal embedding is projected to match the token embedding dimensions required by each LLM. All models are permissively licensed for academic research, and training takes approximately 8 GPU hours per LLM training cycle. During inference, we use a batch size of 1, with generation parameters such as top_k, top_p, and temperature set to their default values for each LLM.
5.3 Evaluation
We use standard NLG metrics such as BLEU, METEOR, ROUGE Sharma et al. (2017), and BERTScore Zhang et al. (2019). Furthermore, we assess the quality of the generation using GPT-4, following Liu et al. (2024b).GPT-4 measures two aspects: fluency, for grammar , and adequacy, for accuracy in conveying meaning, both rated on a scale of 1-5, with 5 denoting the highest quality.
| LLM | ROUGE-N | ROUGE-L | BLEU-N | MET- | BERT | GPT-4 | GPT-4 | ||
| N=1 | N=2 | N=1 | N=4 | EOR | Score | Flu. | Ade. | ||